A ambulance is mooving with a speed v=30ms-1 and a car moving towards the ambulance with the speed of v=15ms-1.
The siran of the ambulance emit sound waves at a frequency 800 HZ.at what frequency will a person in the car hear this sound wave.
1
Expert's answer
2021-10-15T10:24:00-0400

The observed frequency for the person in the car is


"f=\frac{c+v_r}{c-v_s}f_0=\frac{343+15}{343-30}800=915\text{ Hz}."